OpenStudy (anonymous):

Given that x⊗y=x⋅y-3 and that x⊙y=5⋅(x-y) Evaluate this expression:(5⊗(4⊗3))⊙(-5)

OpenStudy (asadkarim7):

its of all solve inner bracket therefore x = 4 , y = 3 4.3-3 = 9 then it would become 5⊗9 that means x= 5 and y= 9 of the same equation 5.9-3= 42 it becomes 42⊙-5 x= 42 , y = -5 5 ( 42-(-5)) 5.47= 235

OpenStudy (anonymous):

correct